# Fast, Stable and Efficient Approximation of Multi-parameter Persistence Modules with MMA

### v3

@@ 2 Background
In this section, we recall the basics of multi-parameter persistence modules. This section only contains the necessary background and notations, and can be skipped if the reader is already familiar with persistence theory. A more complete treatment of persistence modules can be found in [CdSGO16, DW22, Oud15].
@@
In their most general form, multi-parameter persistence modules are nothing but $k$-vector spaces (where $k$ denotes a field) indexed by $\mathbb{R}^{n}$ and connected by linear maps.
Definition 2.1 (Multi-parameter persistence module).
An ($n$-)multi-parameter persistence module ${\mathbb{M}}$ is a family of vector spaces indexed over $\mathbb{R}^{n}$: ${\mathbb{M}}=\{M_{x}\}_{x\in\mathbb{R}^{n}}$, equipped with linear transformations $\{\varphi_{x}^{y}:M_{x}\rightarrow M_{y}\}_{x,y\in\mathbb{R}^{n},x\leq_{n}y}$, that are called the transition maps of ${\mathbb{M}}$, and that satisfy $\varphi_{x}^{z}=\varphi_{y}^{z}\circ\varphi_{x}^{y}$ for any $x\leq_{n}y\leq_{n}z$.
@@
In this article, all multi-parameter persistence modules come from applying the homology functor $H_{*}$ on a multi-parameter filtration of a simplicial complex $S$, that is, on a family $\{S_{x}\}_{x\in\mathbb{R}^{n}}$ of subsets of $S$ indexed over $\mathbb{R}^{n}$ such that $x\leq_{n}y\Rightarrow S_{x}\subseteq S_{y}$. In other words, we study modules of the form ${\mathbb{M}}:=\{H_{*}(S_{x})\}_{x\in\mathbb{R}^{n}}$, where the linear maps $H_{*}(S_{x})\to H_{*}(S_{y})$ are induced by the canonical inclusions $S_{x}\subseteq S_{y}$ (when $x\leq_{n}y$). There are many interesting multi-parameter filtrations in data science; one of the most common one (with $n=2$) comes from filtering by feature scale and density. This allows to detect the topological structures (encoded in the homology groups) of point clouds in the face of noise and outliers [BL20, CB20].
Definition 2.2.
The direct sum of two multi-parameter persistence modules ${\mathbb{M}}$ and ${\mathbb{M}}^{\prime}$, written as ${\mathbb{M}}\oplus{\mathbb{M}}^{\prime}$, is the module ${\mathbb{M}}^{\prime\prime}$ with vector spaces $\{M^{\prime\prime}_{x}\}_{x\in\mathbb{R}^{n}}$ and transition maps (ϕ′′)superscriptsubscriptsuperscriptitalic-$(\phi^{\prime\prime})_{\cdot}^{\cdot}$, defined as $M^{\prime\prime}_{x}=M_{x}\oplus M^{\prime}_{x}$ for all $x\in\mathbb{R}^{n}$, and $(\phi^{\prime\prime})_{\cdot}^{\cdot}=\phi_{\cdot}^{\cdot}\oplus(\phi^{\prime})_{\cdot}^{\cdot}$, where $\{M_{x}\}_{x\in\mathbb{R}^{n}}$ (resp. $\{M^{\prime}_{x}\}_{x\in\mathbb{R}^{n}}$) and ϕsuperscriptsubscriptitalic-$\phi_{\cdot}^{\cdot}$ (resp. (ϕ)superscriptsubscriptsuperscriptitalic-$(\phi^{\prime})_{\cdot}^{\cdot}$) are the vector spaces and transition maps of ${\mathbb{M}}$ (resp. ${\mathbb{M}}^{\prime}$) respectively.
A multi-parameter persistence module ${\mathbb{M}}$ such that there are no modules $A$ and $B$ such that ${\mathbb{M}}\simeq A\oplus B$ is called indecomposable.
Note that while multi-parameter persistence modules can always be decomposed into indecomposable summands (see [DX22] for a corresponding algorithm), these summands can be arbitrarily complicated, and the resulting decomposition cannot really be used as an intuitive and simple invariant of the module.
@@
Multi-parameter persistence modules can be compared with the standard interleaving distance [Les15].
Definition 2.3 (Interleaving distance).
@@
The main property of this distance is that it is stable for multi-parameter filtrations that are obtained from the sublevel sets of functions. More precisely, given two continuous functions $f,g:S\to\mathbb{R}^{n}$ defined on a simplicial complex $S$, and such that $\tau\subseteq\sigma\Rightarrow f(\tau)\leq_{n}f(\sigma)$ for any $\tau,\sigma\in S$ (and similarly for $g$), let ${\mathbb{M}}(f),{\mathbb{M}}(g)$ denote the multi-parameter persistence modules obtained from the corresponding multi-parameter filtrations $\{S^{f}_{x}:=\{\sigma\in S:f(\sigma)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$ and $\{S_{x}^{g}:=\{\sigma\in S:g(\sigma)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$. Then, one has [Les15, Theorem 5.3]:
Another usual distance is the bottleneck distance [BL18, Section 2.3]. Intuitively, it relies on decompositions of the modules into direct sums of indecomposable summands, and is defined as the largest interleaving distance between summands that are matched under some matching.
Definition 2.4 (Bottleneck distance).
Given two multisets $A$ and $B$, $\mu\colon A\not\to B$ is called a matching if there exist $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $\mu\colon A^{\prime}\to B^{\prime}$ is a bijection. The subset $A^{\prime}:=\mathrm{coim}(\mu)$ (resp. $B^{\prime}:=\mathrm{im}(\mu)$) is called the coimage (resp. image) of $\mu$.
Let ${\mathbb{M}}\cong\bigoplus_{i\in\mathcal{I}}M_{i}$ and ${\mathbb{M}}^{\prime}\cong\bigoplus_{j\in\mathcal{J}}M^{\prime}_{j}$ be two multi-parameter persistence modules. Given $\varepsilon\geq 0$, the modules ${\mathbb{M}}$ and ${\mathbb{M}}^{\prime}$ are $\boldsymbol{\varepsilon}$-matched if there exists a matching $\mu\colon\mathcal{I}\not\to\mathcal{J}$ such that $M_{i}$ and $M^{\prime}_{\mu(i)}$ are $\boldsymbol{\varepsilon}$-interleaved for all $i\in\mathrm{coim}(\mu)$, and $M_{i}$ (resp. $M^{\prime}_{j}$) is $\boldsymbol{\varepsilon}$-interleaved with the null module 0 for all $i\in\mathcal{I}\backslash\mathrm{coim}(\mu)$ (resp. $j\in\mathcal{J}\backslash\mathrm{im}(\mu)$).
[…]
Since a matching between the decompositions of two multi-parameter persistence modules induces an interleaving between the modules themselves, it follows that ${d_{\rm I}}\leq{d_{\rm B}}$. Note also that ${d_{\rm B}}$ can actually be arbitrarily larger than ${d_{\rm I}}$, as showcased in [BL18, Section 9].
@@
Our approximation result presented in Section 4.2 is stated for modules that can be decomposed into so-called interval modules. Hence, in this section, we define such interval modules. Intuitively, they are modules that are trivial, except on a subset of $\mathbb{R}^{n}$ called an interval.
Definition 2.5 (Interval).
@@
Definition 2.6 (Indicator module, Interval module).
@@
A multi-parameter persistence module ${\mathbb{M}}$ is an indicator module if there exists a set $S\subseteq\mathbb{R}^{n}$, called the support of ${\mathbb{M}}$ and denoted by ${\rm supp}({\mathbb{M}})$, such that:
$\forall x\in\mathbb{R}^{n},\,M_{x}=\begin{cases}k\text{ if }x\in S\\ {\bf 0}\text{ otherwise}\end{cases}\ \ \ {\rm and}\ \ \ \,\forall x,y\in\mathbb{R}^{n},\,\varphi_{x}^{y}=\begin{cases}\mathrm{id}_{k\rightarrow k}\text{ if }x\leq_{n}y\in S\\ 0\text{ otherwise}\end{cases}$
where $\varphi_{\cdot}^{\cdot}$ are the transition maps of ${\mathbb{M}}$. If $S$ is an interval, ${\mathbb{M}}$ is called an interval module.
@@
Finally, we define a specific type of interval modules, those whose support is equal to a union of rectangles. We call these modules discretely presented—the stable candidate decompositions computed by our algorithm MMA in Section 4 are actually made up of such modules.
Definition 2.7 (Discretely presented interval module).
@@
An interval module $I$ is discretely presented if its support is a locally finite union of rectangles in ¯$\overline{\mathbb{R}}^{n}$, and whose boundary is an $(n-1)$-submanifold of $\mathbb{R}^{n}$. More precisely, there exist two locally finite families of points, the birth and death critical points of $I$, denoted by $C_{B}(I)$ and $C_{D}(I)$ respectively, such that:
$I:=\mathrm{\mathbf{Int}}\left(\bigcup_{c\in C_{B}(I)}\bigcup_{c^{\prime}\in C_{D}(I)}R_{c,c^{\prime}}\right),$ (2)
where $R_{c,c^{\prime}}:=\{x\in\mathbb{R}^{n}:c\leq_{n}x\leq_{n}c^{\prime}\}$ is the rectangle defined by corners $c$ and $c^{\prime}$, and $\mathrm{\mathbf{Int}}\left(S\right)$ (where $S$ is an interval) denotes the interval module whose support is $S$.
@@
Note that when only one filtration is given, single-parameter persistence modules always decompose into interval modules: ${\mathbb{M}}\cong\bigoplus_{i\in\mathcal{I}}\mathrm{\mathbf{Int}}\left([b_{i},d_{i}]\right)$ [CdSGO16, Theorem 2.8]. In that case, they are frequently represented as the collection of the interval supports (in $\mathbb{R}$) of their summands, also called persistence barcode ${\mathcal{B}}({\mathbb{M}}):=\{[b_{i},d_{i}]\}_{i\in\mathcal{I}}$.
@@
The fibered barcode [LW15] is a centerpiece of our approximation scheme, and is defined, given a multi-parameter persistence module ${\mathbb{M}}$, as a map that takes as input a line (or segment) $l$ in $\mathbb{R}^{n}$, and outputs the persistence barcode associated to the single-parameter persistence module obtained by restricting ${\mathbb{M}}$ along $l$. Hence, in the following, we formalize the intersections between multi-parameter persistence modules and lines in $\mathbb{R}^{n}$.
Definition 2.8 (Fibered Barcode).
Let ${\mathbb{M}}$ be a pointwise finite-dimensional multi-parameter persistence module. Given a discrete family of diagonal lines $L\subseteq\mathcal{L}$ (i.e., lines with direction vector $(1,\dots,1)\in\mathbb{R}^{n}$), we let the $L$-fibered barcode (or fibered barcode for short when $L$ is clear) be the family of barcodes associated to restrictions of the module to lines in $L$, i.e., ${\mathcal{F}}{\mathcal{B}}({\mathbb{M}})_{L}=\{{\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}}):l\in L\}$.
Remark 2.9.
Recall from single-parameter persistence theory that the support of the restriction of ${\mathbb{M}}$ to $l$ is a set of bars called persistence barcode: ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})={\rm supp}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})=\{[b_{i},d_{i}]\}_{i\in\mathcal{I}(l)}$, where $\mathcal{I}(l)$ is an index set that depends on ${\mathbb{M}}$ and $l$. Moreover, when ${\mathbb{M}}=\bigoplus_{i\in\mathcal{I}}I_{i}$ is decomposable into interval modules, there are as many bars in the barcode as there are interval summands intersecting the line ${\mathcal{B}}({\left.\kern-1.2pt{\mathbb{M}}\vphantom{\big{|}}\right|_{l}})=\{{\rm supp}({\left.\kern-1.2ptI_{i}\vphantom{\big{|}}\right|_{l}})\}_{i\in\mathcal{I}}$.
@@
Definition 2.10 (Birthpoint, Deathpoint).
Given a point $x\in\mathbb{R}^{n}$ and an interval module $I$, we call $b_{x}^{I}:=\min\left\{x+\boldsymbol{\delta}\,:\,\delta\in\mathbb{R}\right\}$ (resp. $d_{x}^{I}=\max\left\{x+\boldsymbol{\delta}\,:\,\delta\in\mathbb{R}\right\}$) the birthpoint (resp. deathpoint) associated to $x$ and $I$, where $\boldsymbol{\delta}=(\delta,\dots,\delta)\in\mathbb{R}^{n}$.
Similarly, given a diagonal line $l\subseteq\mathbb{R}^{n}$, we define the birthpoint (resp. deathpoint) associated to $l$ and $I$ as $b_{l}^{I}:=b_{x}^{I}$ (resp. $d_{l}^{I}:=d_{x}^{I}$) for any $x\in l$. See Figure 2
Remark 2.11.

### v4

@@ 2 Background
In this section, we recall the basics of multi-parameter persistent homology and persistence modules. This section only contains the necessary background and notations, and can be skipped if the reader is already familiar with persistence theory. A more complete treatment of persistence modules can be found in (24, 31, 54, 13).
Notations.
We first introduce a few notations: we let $(e_{1},\dots,e_{n})$ be the canonical basis of $\mathbb{R}^{n}$, and, given a set $A\subseteq\mathbb{R}^{n}$, we let ${\rm conv}(A)$ denote the convex hull of $A$. Moreover, given a hyperplane $H\subseteq\mathbb{R}^{n}$ and its two associated vectors $a_{H},b_{H}\in\mathbb{R}^{n}$ which satisfy $H=b_{H}+\{x\in\mathbb{R}^{n}:\langle x,a_{H}\rangle=0\}$, we call $a_{H}$ the codirection of $H$. When $a_{H}$ is a vector in the canonical basis of $\mathbb{R}^{n}$, i.e., there exists $i\in\llbracket 1,n\rrbracket$ such that $a_{H}=e_{i}$, we slightly abuse notation and also call $i$ the codirection of $H$. Finally, given a point $x\in\mathbb{R}^{n}$, we let $\left<x\right$< (resp. $\left>x\right$>) denote the upset (resp. downset) of $x$: $\left<x\right<:=\left\{y\in\mathbb{R}^{n}:y\geq_{n}x\right\}$ (resp. $\left>x\right>:=\left\{y\in\mathbb{R}^{n}:y\leq_{n}x\right\}$).
@@
In their most general form, multi-parameter persistence modules (19) are nothing but $k$-vector spaces (where $k$ denotes a field) indexed by $\mathbb{R}^{n}$ and connected by linear maps.
Definition 1 (Multi-parameter persistence module).
An ($n$-)multi-parameter persistence module $\mathbb{M}$ is a family of vector spaces indexed over $\mathbb{R}^{n}$: $\mathbb{M}=\{M_{x}\}_{x\in\mathbb{R}^{n}}$, equipped with linear transformations $\{\varphi_{x}^{y}:M_{x}\rightarrow M_{y}\}_{x,y\in\mathbb{R}^{n},x\leq_{n}y}$, that are called the transition maps of $\mathbb{M}$, and that satisfy $\varphi_{x}^{z}=\varphi_{y}^{z}\circ\varphi_{x}^{y}$ and $\varphi^{x}_{x}=\mathrm{id}$ for any $x\leq_{n}y\leq_{n}z$. We sometimes let $\mathbb{M}(x\leq_{n}y):=\varphi_{x}^{y}$ denote these transition maps.
@@
Multi-parameter persistence modules are often assumed to satisfy finiteness assumptions, such as being pointwise finite dimensional (Definition 2) or finitely presentable (Definition 9). See (13) for more details.
Definition 2 (Pointwise finite dimensional module).
Let $\mathbb{M}$ be an $n$-parameter persistence module. We say that $\mathbb{M}$ is pointwise finite dimensional (p.f.d.), if for any $x\in\mathbb{R}^{n}$, one has $\dim\mathbb{M}_{x}<\infty$.
In this article, all multi-parameter persistence modules come from applying the homology functor $H_{*}$ on a multi-parameter filtration of a topological space $S$, that is, on a family $\{S_{x}\}_{x\in\mathbb{R}^{n}}$ of subsets of $S$ indexed over $\mathbb{R}^{n}$ such that $x\leq_{n}y\Rightarrow S_{x}\subseteq S_{y}$. In other words, we study modules of the form $\mathbb{M}:=\{H_{*}(S_{x})\}_{x\in\mathbb{R}^{n}}$, where the linear maps $H_{*}(S_{x})\to H_{*}(S_{y})$ are induced by the canonical inclusions $S_{x}\subseteq S_{y}$ (when $x\leq_{n}y$). There are many interesting multi-parameter filtrations in data science; one of the most common one (with $n=2$) comes from filtering by feature scale and density. This allows to detect the topological structures (encoded in the homology groups) of point clouds in the face of noise and outliers (9, 20).
Definition 3.
The direct sum of two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$, written as $\mathbb{M}\oplus\mathbb{M}^{\prime}$, is the module $\mathbb{M}^{\prime\prime}$ with vector spaces $\{M^{\prime\prime}_{x}\}_{x\in\mathbb{R}^{n}}$ and transition maps $(\varphi^{\prime\prime})_{\cdot}^{\cdot}$, defined as $M^{\prime\prime}_{x}=M_{x}\oplus M^{\prime}_{x}$ for all $x\in\mathbb{R}^{n}$, and $(\varphi^{\prime\prime})_{\cdot}^{\cdot}=\varphi_{\cdot}^{\cdot}\oplus(\varphi^{\prime})_{\cdot}^{\cdot}$, where $\{M_{x}\}_{x\in\mathbb{R}^{n}}$ (resp. $\{M^{\prime}_{x}\}_{x\in\mathbb{R}^{n}}$) and $\varphi_{\cdot}^{\cdot}$ (resp. $(\varphi^{\prime})_{\cdot}^{\cdot}$) are the vector spaces and transition maps of $\mathbb{M}$ (resp. $\mathbb{M}^{\prime}$) respectively.
A multi-parameter persistence module $\mathbb{M}$ such that there are no non-trivial modules $A$ and $B$ such that $\mathbb{M}\simeq A\oplus B$ is called indecomposable.
Note that while multi-parameter persistence modules can always be decomposed into indecomposable summands if they are p.f.d. (see (34, 32) for corresponding algorithms), these summands can be arbitrarily complicated, and the resulting decomposition cannot really be used as an intuitive and simple invariant of the module.
@@
Multi-parameter persistence modules can be compared with the standard interleaving distance (45).
Definition 4 (Interleaving distance).
@@
The interleaving distance between two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$ is then defined as ${d_{\rm I}}(\mathbb{M},\mathbb{M}^{\prime}):=\inf\left\{\varepsilon\geq 0:\mathbb{M}\text{ and }\mathbb{M}^{\prime}\text{ are }\boldsymbol{\varepsilon}\text{-interleaved}\right\}$.
The main property of this distance is that it is stable for multi-parameter filtrations that are obtained from the sublevel sets of functions. More precisely, given two continuous functions $f,g:S\to\mathbb{R}^{n}$ defined on a topological space $S$ let $\mathbb{M}(f),\mathbb{M}(g)$ denote the multi-parameter persistence modules obtained from the corresponding multi-parameter filtrations $\{S^{f}_{x}:=\{s\in S:f(s)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$ and $\{S_{x}^{g}:=\{s\in S:g(s)\leq_{n}x\}\}_{x\in\mathbb{R}^{n}}$. Then, one has (45, Theorem 5.3):
Another usual distance is the bottleneck distance (12, Section 2.3). Intuitively, it relies on decompositions of the modules into direct sums of indecomposable summands, and is defined as the largest interleaving distance between summands that are matched under some matching.
Definition 5 (Bottleneck distance).
Given two multisets $A$ and $B$, $\mu\colon A\not\to B$ is called a partial bijection if there exist $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ such that $\mu\colon A^{\prime}\to B^{\prime}$ is a bijection. The subset $A^{\prime}:=\mathrm{coim}(\mu)$ (resp. $B^{\prime}:=\mathrm{im}(\mu)$) is called the coimage (resp. image) of $\mu$.
@@
The bottleneck distance between two multi-parameter persistence modules $\mathbb{M}$ and $\mathbb{M}^{\prime}$ is then defined as ${d_{\rm B}}(\mathbb{M},\mathbb{M}^{\prime}):=\inf\left\{\varepsilon\geq 0\,:\,\mathbb{M}\text{ and }\mathbb{M}^{\prime}\text{ are }\boldsymbol{\varepsilon}\text{-matched}\right\}.\\$
Since a matching between the decompositions of two multi-parameter persistence modules induces an interleaving between the modules themselves, it follows that ${d_{\rm I}}\leq{d_{\rm B}}$. Note also that $d_{\rm B}$} can actually be arbitrarily larger than $d_{\rm I}$}, as showcased in (12, Section 9).
@@
Now, we define a particular subfamily of multi-parameter persistence modules, the so-called interval modules. Intuitively, they are modules that are trivial, except on a subset of $\mathbb{R}^{n}$ called an interval.
Definition 6 (Interval).
@@
Definition 7 (Interval module, indicator module).
@@
An interval module $\mathbb{I}$ is a multi-parameter persistence module such that:
  1. 1.
    $\mathbb{I}$ is a thin module, i.e., $\forall x\in\mathbb{R}^{n},\,{\rm dim}(\mathbb{I}_{x})\leq 1$,
  2. 2.
    whose support ${\rm supp}(\mathbb{I}):=\{x\in\mathbb{R}^{n}:{\rm dim}(\mathbb{I}_{x})=1\}$ is an interval of $\mathbb{R}^{n}$,
  3. 3.
    and whose transition maps are identity maps, i.e., $\forall x\leq_{n}y\in{\rm supp}(\mathbb{I}),\,\mathbb{I}(x\leq y)=\mathrm{id}$.
Moreover, given an interval $I$, we let $k^{I}$ denote the corresponding interval module with support $I$. Finally, a module $\mathbb{I}$ that is a direct sum of interval modules $\mathbb{I}=k^{I_{1}}\oplus\dots\oplus k^{I_{m}}$, $m\in\mathbb{N}^{*}$, whose supports have empty pairwise (closed) intersections, i.e., such that $I_{p}\cap I_{q}=\varnothing$ for all $1\leq p,q\leq m$, $p\neq q$, is called an indicator module, and denoted by $k^{U}$, where $U:=\cup_{i=1}^{m}I_{i}$ is the union of their supports.
@@
Finally, we define a specific type of interval modules, those whose support is equal to a union of rectangles. We call these modules discretely presented—the candidate decompositions computed by our algorithm MMA (Section 3) are actually made up of such modules.
Definition 8 (Discretely presented interval module).
@@
An interval module $\mathbb{I}=k^{I}$ is discretely presented if its support $I$ is a locally finite union of rectangles in $\overline{\mathbb{R}}^{n}$, and whose boundary is an (n1)(n-1)-submanifold of $\mathbb{R}^{n}$. More precisely, there exist two locally finite families of points, the birth and death critical points of $I$, denoted by $C_{B}(I)$ and $C_{D}(I)$ respectively, such that:
$I:=\bigcup_{c\in C_{B}(I)}\bigcup_{c^{\prime}\in C_{D}(I)}R_{c,c^{\prime}},$ (2)
where $R_{c,c^{\prime}}:=\{x\in(\mathbb{R}\cup\{\pm\infty\})^{n}:c\leq_{n}x{{<}_{n}}c^{\prime}\}$ is the rectangle with corners $c$ and $c^{\prime}$.
Discretely presented interval modules can be obtained through stronger assumptions, such as being a finitely presentable interval module. See (13) for more details.
Definition 9 (Finitely presentable persistence module.).
Let $\mathbb{M}$ be an $n$-parameter persistence module. One says that $\mathbb{M}$ is finitely presentable (f.p.) if it is isomorphic to the cokernel of a persistence morphism $f\colon R\to G$, where $G,R$ are interval decomposable modules of the form:
$G=\bigoplus_{1\leq i\leq n_{G}}k^{\left<x_{i}\right<}\quad\textnormal{ and }\quad R=\bigoplus_{1\leq i\leq n_{R}}k^{\left<y_{i}\right<},$ (3)
for some finite sets $\beta_{0}:=(x_{i})_{1\leq i\leq n_{G}}$ and $\beta_{1}:=(y_{i})_{1\leq i\leq n_{R}}$, with $n_{G},n_{r}\in\mathbb{N}$. When $n_{G}$ and $n_{R}$ are minimal, the sets $\beta_{0}(\mathbb{M})$ and $\beta_{1}(\mathbb{M})$ are unique and called the degree 0 and degree 1 graded Betti numbers of $\mathbb{M}$, respectively.
The graded Betti numbers characterize the position of topological events in a multi-parameter filtration. In particular, f.p. modules can be restricted to large enough compact sets without loosing information, if such sets contain the graded Betti numbers, see Remark 1. Next, we define restrictions of modules.
Definition 10 (Restrictions and slices).
If $Q\subseteq\mathbb{R}^{n}$, then an $n$-parameter persistence module $\mathbb{M}=\{M_{x}\}_{x\in\mathbb{R}^{n}}$ induces a persistence module $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}$} indexed over $Q$, defined by:
$\forall x\leq_{n}y\in Q,\quad\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}}\right)_{x}:=M_{x}\textnormal{ and }{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{Q}}(x\leq_{n}y):=\mathbb{M}(x\leq_{n}y).$
In particular, if $\mathbb{M}$ is an $n$-parameter persistence module, then given a line $l\subseteq\mathbb{R}^{n}$ with positive slope, the persistence module $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}$} can be seen as a usual 1-parameter persistence module up to some parametrization of $l$.
Example 1 (Restriction to a line).
[…]
$\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)_{t}:=M_{at+b}\quad\textnormal{and }\quad\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)(s\leq t):=\mathbb{M}(as+b\leq at+b),$ (4)
is a $1$-parameter module, called the restriction, or slice, of $\mathbb{M}$ along $l$.
The opposite operation is called a (left) Kan extension, that we only define in our setup.
Definition 11 (Kan Extension).
Let $A=A_{1}\times\cdots\times A_{n}\subseteq\mathbb{R}^{n}$ be a poset of $\mathbb{R}^{n}$ obtained as a product of subsets of $\mathbb{R}$, and $\mathbb{M}=\{M_{x}\}_{x\in A}$ be a multi-parameter persistence module indexed over $A$. The (left) Kan extensions of $\mathbb{M}$ is the $n$-parameter persistence module (indexed over $\mathbb{R}^{n}$) defined as follows: for any $x\leq_{n}y\in\mathbb{R}^{n}$,
$\left(\mathrm{Lan}_{n}\mathbb{M}\right)_{x}=M_{\lfloor x\rfloor_{A}}\quad\textnormal{and}\quad\mathrm{Lan}_{n}\mathbb{M}(x\leq_{n}y)=\mathbb{M}(\lfloor x\rfloor_{A}\leq_{n}\lfloor y\rfloor_{A}),$ (5)
[…]
Remark 1.
If the graded Betti numbers of $\mathbb{M}$ are included in a rectangle subset $K=[a_{1},b_{1}]\times\cdots\times[a_{n},b_{n}]$ of $\mathbb{R}^{n}$, then the restriction $\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}$} contains as much information as the original module $\mathbb{M}$; more formally, we have $\mathrm{Lan}_{n}{\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{K}}\simeq\mathbb{M}$. Notice that, in practice, finding such a compact set $K$ for a given persistence module $\mathbb{M}$ does not require computing a minimal presentation of $\mathbb{M}$ (Definition 9). Indeed, if $\mathbb{M}$ is obtained from applying the homology functor on a finite multi-parameter filtration $F$, i.e., $\mathbb{M}\simeq H_{*}(F)$, then any rectangle $R\subseteq\mathbb{R}^{n}$ containing the filtration values of the simplices of $F$ will also contain the graded Betti numbers.
@@
Note that when only one filtration is given, single-parameter p.f.d. persistence modules (Definition 2) always decompose into interval modules: $\mathbb{M}\cong\bigoplus_{i\in\mathcal{I}}k^{[b_{i},d_{i})}$ (24, Theorem 2.8). In that case, they are frequently represented as the collection of the supports (in $\mathbb{R}$) of their summands, also called persistence barcode ${\mathcal{B}}\left(\mathbb{M}\right):=\{[b_{i},d_{i})\}_{i\in\mathcal{I}}$.
Boundaries and facets of intervals
Now, we recall the definition of upper- and lower-boundaries of interval modules, as well as their so-called facets, which are convenient characterizations of the interval supports.
Definition 12 (Upper- and lower-boundaries).
Given an interval $I\subset\mathbb{R}^{n}$, its upper-boundary $U[I]$ and lower-boundary $L[I]$ are defined as:
$L[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y<_{n}x\Rightarrow y\not\in I\right\},\ \ U[I]:=\left\{x\in\bar{I}:\forall y\in\mathbb{R}^{n},\,y>_{n}x\Rightarrow y\not\in I\right\}.$
Moreover, the boundary of I{I} can be decomposed with $\partial{I}=L[I]\cup U[I]$. See Figure 3 for an illustration.
Figure 3: Lower- and upper-boundaries of an interval in $\mathbb{R}^{2}$ (Definition 12); and birthpoints and deathpoints $b_{x}^{I}$ and $d_{x}^{I}$ (Definition 15) of a point $x\in\mathbb{R}^{2}$.
When interval modules are discretely presented (see Definition 8), their lower- and upper-boundaries are made of flat parts, which are the faces of the corresponding rectangles forming the interval. Hence, we call facets the subsets of the lower- and upper-boundaries that are included in some hyperplanes of $\mathbb{R}^{n}$.
Definition 13 (Facet).
A lower (resp. upper) facet of an interval $I\subset\mathbb{R}^{n}$ is an (n1)(n-1)-submanifold of $\partial{\rm supp}(I)$ written as $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap L[I]$ (resp. $\left\{x\in\mathbb{R}^{n}\,:\,x_{i}=c\right\}\cap U[I]$) for some $c\in\mathbb{R}$ and some dimension $i\in\llbracket 1,n\rrbracket$ that is called the facet codirection. In particular, the upper- and lower-boundaries of a discretely presented interval module is a locally finite union of facets.
@@
The fibered barcode (23, 47) is a centerpiece of our MMA algorithm, and is defined, given a p.f.d. multi-parameter persistence module $\mathbb{M}$, as a map that takes as input a line (or segment) $l$ in $\mathbb{R}^{n}$, and outputs the persistence barcode associated to the single-parameter persistence module obtained by restricting $\mathbb{M}$ along $l$. We formalize these concepts in the next definition.
Definition 14 ( Diagonal fibered barcode).
Let $\mathbb{M}$ be a p.f.d. multi-parameter persistence module. Given a set $L$ of diagonal lines in $\mathbb{R}^{n}$ (i.e., lines with direction vector $1,\dots,1]\in\mathbb{R}^{n}$), we let the $L$-fibered barcode (or fibered barcode for short when $L$ is clear) be the family of barcodes associated to restrictions of the module along lines in $L$, i.e., ${\mathcal{F}}{\mathcal{B}}(\mathbb{M})_{L}={\left({\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)\right)_{l\in L}}$.
Remark 2.
Recall from (24, Theorem 2.8) that ${\rm supp}({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})$ is a multiset of bars called persistence barcode: ${\rm supp}({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}})={\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)=\{[b_{i},d_{i})\}_{i\in\mathcal{I}(l)}$, where $\mathcal{I}(l)$ is an index set that depends on $\mathbb{M}$ and $l$. Moreover, when $\mathbb{M}=\bigoplus_{i\in\mathcal{I}}k^{I_{i}}$ is decomposable into interval modules, there are as many bars in the barcode as there are interval summands intersecting the line $l$: $|{\mathcal{B}}\left({\left.\kern-1.2pt\mathbb{M}\vphantom{\big|}\right|_{l}}\right)|=|\{I_{i}:I_{i}\cap l\neq\varnothing\}|$.
@@
Definition 15 (Birthpoint, Deathpoint).
Given a positive line $l$ (that is, a line whose direction vector $u$ is in $\mathbb{R}^{n}_{+}\setminus\left\{0\right\}$) and an interval $I\subseteq\mathbb{R}^{n}$, the birthpoint (resp. deathpoint) of $I$ along $l$ is:
$b^{I}_{l}:=\inf\,l\cap I,\quad\text{resp. }d^{I}_{l}:=\sup\,l\cap I.$
If the direction $u\in\mathbb{R}_{+}^{n}$ of the line is given by the context, and $x\in\mathbb{R}^{n}$, we will also let $b^{I}_{x}:=b_{l_{x}}^{I}$ (resp. $d^{I}_{x}:=d_{l_{x}}^{I}$) denote the birthpoint (resp. deathpoint) associated to $I$ and $x$, where $l_{x}$ is the line crossing $x$, with direction vector $u$. See Figure 3.
Remark 3 (Slicing an interval decomposable module).